# OpenXAI: Towards a Transparent Evaluation of Model Explanations

### v1

@@ Abstract
While several types of post hoc explanation methods (e.g., feature attribution methods) have been proposed in recent literature, there is little to no work on systematically benchmarking these methods in an efficient and transparent manner. Here, we introduce OpenXAI, a comprehensive and extensible open-source framework for evaluating and benchmarking post hoc explanation methods. OpenXAI comprises of the following key components: (i) a flexible synthetic data generator and a collection of diverse real-world datasets, pre-trained models, and state-of-the-art feature attribution methods, (ii) open-source implementations of twenty-two quantitative metrics for evaluating faithfulness, stability (robustness), and fairness of explanation methods, and (iii) the first ever public XAI leaderboards to benchmark explanations. OpenXAI is easily extensible, as users can readily evaluate custom explanation methods and incorporate them into our leaderboards. Overall, OpenXAI provides an automated end-to-end pipeline that not only simplifies and standardizes the evaluation of post hoc explanation methods, but also promotes transparency and reproducibility in benchmarking these methods. OpenXAI datasets and data loaders, implementations of state-of-the-art explanation methods and evaluation metrics, as well as leaderboards are publicly available at https://open-xai.github.io/.
@@ 1 Introduction
As predictive models are increasingly deployed in critical domains (e.g., healthcare, law, and finance), there has been a growing emphasis on explaining the predictions of these models to decision makers (e.g. doctors, judges) so that they can understand the rationale behind model predictions, and determine if and when to rely on these predictions. To this end, various techniques have been proposed in recent literature to generate post hoc explanations of individual predictions made by complex ML models. Several of such local explanation methods output the influence of each of the features on the model’s prediction, and are therefore referred to as feature attribution methods. Due to their generality, feature attribution methods are increasingly being utilized to explain complex models in medicine, finance, law, and science [15, 25, 65]. Thus, it is critical to ensure that the explanations generated by these methods are reliable so that relevant stakeholders and decision makers are provided with credible information about the underlying models [2].
Prior works have studied several notions of explanation reliability such as faithfulness (or fidelity) [68, 41, 22], stability (or robustness) [3, 1], and fairness [12, 5], and proposed metrics for quantifying these notions. Many of these works also demonstrated through small-scale experiments or qualitative analysis that certain explanation methods are not effective w.r.t. specific notions of reliability. For instance, Alvarez-Melis and Jaakkola [3] visualized gradient-based explanations [56, 55, 59, 60] for MNIST images and showed that they are not robust to small input perturbations. However, it is unclear if such findings generalize beyond the settings studied. More broadly, one of the biggest open questions which has far-reaching implications for the progress of explainable AI (XAI) research is: which explanation methods are effective w.r.t. what notions of reliability and under what conditions? [34]. A first step towards answering this question involves systematically benchmarking explanation methods in a reproducible and transparent manner. However, the increasing diversity of explanation methods, and the plethora of evaluation settings and metrics outlined in existing research without standardized open-source implementations make it rather challenging to carry out such benchmarking efforts.
@@
    Our OpenXAI framework currently provides open-source implementations and ready-to-use API interfaces for seven state-of-the-art feature attribution methods (LIME, SHAP, Vanilla Gradients, Gradient x Input, SmoothGrad, and Integrated Gradients), and twenty-two quantitative metrics to evaluate the faithfulness, stability, and fairness of feature attribution methods. In addition, it includes a comprehensive collection of five real-world datasets spanning diverse real-world domains, and twelve different pre-trained models. OpenXAI also introduces a novel and flexible synthetic data generator to synthesize datasets of varying sizes, complexity, and dimensionality which facilitate the construction of reliable ground truth explanations (Section 2).
@@
    As part of our OpenXAI framework, we also develop the first ever public XAI leaderboards (shown in Figure 2) to promote transparency, and to allow users to easily compare the performance of multiple explanation methods across a wide variety of evaluation metrics, predictive models, and datasets.
@@
    OpenXAI framework is easily extensible i.e., researchers and practitioners can readily incorporate custom explanation methods, datasets, predictive models, and evaluation metrics into our framework and leaderboards (Section 2).
  5. 5.
    Lastly, using our proposed OpenXAI framework, we perform rigorous empirical benchmarking of the aforementioned state-of-the-art feature attribution methods to determine which methods are effective w.r.t. what notions of reliability across a wide variety of datasets and predictive models (Section 3).
@@
Related Work. Our work builds on the vast literature in explainable AI. Here, we discuss closely related works and their connections to our benchmark. A more detailed discussion of the related work is included in the Appendix.
@@
Evaluation Metrics for Post hoc Explanations: Prior research has studied several notions of explanation reliability, namely, faithfulness (or fidelity), stability (or robustness), and fairness [41, 68, 3, 12]. While the faithfulness notion captures how faithfully a given explanation captures the true behavior of the underlying model [68, 41, 22], stability ensures that explanations do not change drastically with small perturbations to the input [19, 3]. The fairness notion, on the other hand, ensures that there are no group-based disparities in the faithfulness or stability of explanations [12]. To this end, prior works [41, 58, 68, 3, 12, 22] proposed various evaluation metrics to quantify the aforementioned notions. For instance, Petsiuk et al. [48] measured the change in the probability of the predicted class when important features (as identified by an explanation) are deleted from or introduced into the data instance. A sharp change in the probability implies a high degree of explanation faithfulness. Alvarez-Melis and Jaakkola [3] loosely quantified stability as the maximum change in the resulting explanations when small perturbations are made to a given instance. Dai et al. [12] quantified unfairness of explanations as the difference between the faithfulness (or stability) metric values averaged over instances in the majority and the minority subgroups.
Benchmarking Post hoc Explanations: While there exist several benchmarks in machine learning literature [11, 23, 24, 41], they focus on evaluating the predictive performance of machine learning models or other key model properties such as adversarial robustness. Therefore, these benchmarks cannot be readily used to evaluate the reliability of post hoc explanations as it requires an entirely new set of evaluation metrics, and experimental set up [41]. While some prior works [41, 32] have constructed synthetic datasets with ground truth explanations to benchmark feature attribution methods, this evaluation is limited in its scope as synthetic datasets are not representative of real world data [41]. In addition, models learned using the aforementioned synthetic datasets may not adhere to the ground truth explanations of the underlying data as demonstrated by Faber et al. [16]. Furthermore, while there exist some libraries with open-source implementations of certain quantitative metrics to evaluate the reliability of post hoc explanations [21], they do not focus on systematic and transparent benchmarking of explanation methods. Our OpenXAI framework is complementary to these prior works, but differs in that we consider a much broader set of evaluation metrics and both synthetic and real-world datasets. To the best of our knowledge, OpenXAI is the only benchmark that provides a complete pipeline to evaluate and compare explanations on all the three key dimensions of faithfulness, stability, and fairness, with a dedicated leaderboard to promote transparency and collaboration around evaluations of post hoc explanations.
@@ 2 Overview of OpenXAI Framework
1) Datasets and Predictive Models. The current release of our OpenXAI framework includes a collection of six different synthetic and real-world datasets. While synthetic datasets allow us to construct ground truth explanations which can then be used to evaluate explanations output by state-of-the-art methods, real-world datasets (where it is typically hard to construct ground truth explanations) help us benchmark these methods in a more realistic manner suitable for practical applications [41]. OpenXAI is the first XAI benchmark to comprehensively evaluate explanation methods by leveraging both synthetic and real-world datasets.
Synthetic Datasets: While prior research [41, 32] proposed methods to generate synthetic datasets and corresponding ground truth explanations, they all suffer from a significant drawback as demonstrated by Faber et al. [16] – there is no guarantee that the models trained on these datasets will adhere to the ground truth explanations of the underlying data. This, in turn, implies that evaluating post hoc explanations using the above ground truth explanations would be incorrect since post hoc explanations are supposed to reliably explain the behavior of the underlying model, and not that of the underlying data. To illustrate, let us consider the case where we use aforementioned methods to construct a synthetic dataset with features $A,B,$ $C$, and $D$ such that the ground truth labels only depend on features $A$ and $B$ i.e., the ground truth explanation of the underlying data indicates that features $A$ and $B$ are most important. If we train a model on this data and if features $A$ and $B$ are correlated with $C$ and $D$ respectively, then the resulting model may base its predictions on $C$ and $D$ (and not $A$ and $B$) and still be very accurate. If a post hoc explanation of this model then (correctly) indicates that the most important features of the model are $C$ and $D$, this explanation may be deemed incorrect if we compare it against the ground truth explanation of the underlying data. This problem further exacerbates as we increase the complexity of the ground truth labeling function [16].
@@
Real-world Datasets: In the current release of OpenXAI, we include five real-world datasets that are highly diverse in terms of several key properties. They comprise of data spanning multiple real-world domains (e.g., finance, lending, and criminal justice), varying dataset sizes (e.g., small vs. large-scale), dimensionalities (e.g., low vs. high dimensional), class imbalance ratios, and feature types (e.g., continuous vs. discrete). We focus on tabular data in this release as such data is commonly encountered in real-world applications where explainability is critical [62], and has also been widely studied in XAI literature [41]. Table 1 provides a summary of the real-world datasets currently included in OpenXAI. See Section E.1 in the Appendix for detailed descriptions of individual datasets. While these real-world datasets are primarily drawn from prior research and existing repositories, OpenXAI provides comprehensive data loading and pre-processing capabilities to make these datasets XAI-ready (more details below). We also plan to expand our collection of real-world datasets in the next iteration. Adding a new dataset into our collection is as simple as uploading a .csv file or a .zip folder. Users can also submit requests to incorporate new datasets into the OpenXAI framework by filling a simple form and providing links to the datasets (See Appendix).
Data loaders and pre-trained models: OpenXAI provides a Dataloader class that can be used to load the aforementioned collection of synthetic and real-world datasets as well as any other custom datasets, and ensures that they are XAI-ready. More specifically, this class takes as input the name of an existing OpenXAI dataset or a new dataset (name of the .csv file), and outputs a train set which can then be used to train a predictive model, a test set which can be used to generate local explanations of the trained model, as well as any ground-truth explanations (if and when available). If the dataset already comes with pre-determined train and test splits, this class loads train and test sets from those pre-determined splits. Otherwise, it divides the entire dataset randomly into train (70%) and test (30%) sets. Users can also customize the percentages of train-test splits. The code snippet below shows how to import the Dataloader class and load an existing OpenXAI dataset. from openxai import Dataloader
loader_train, loader_test = Dataloader.return_loaders(data_name=‘german’, download=True)
inputs, labels = iter(loader_test).next()
We also pre-trained two classes of predictive models (e.g., deep neural networks of varying degrees of complexity, logistic regression models etc.) and incorporated them into the OpenXAI framework so that they can be readily used for benchmarking explanation methods. The code snippet below shows how to load OpenXAI’s pre-trained models using our LoadModel class. from openxai import LoadModel
@@
Dataset Size # features Feature types Feature information Balanced
Synthetic Data 5,000 20 continuous synthetic
German Credit [14] 1,000 20 discrete, continuous demographic, personal, financial
HELOC [17] 9,871 23 continuous demographic, financial
COMPAS [27] 18,876 7 discrete, continuous demographic, personal, criminal
Adult Income [66] 48,842 13 discrete, continuous demographic, personal, education/employment, financial
Give Me Some Credit [18] 102,209 10 discrete, continuous demographic, personal, financial
Adding additional pre-trained models into the OpenXAI framework is as simple as uploading a file with details about model architecture and parameters in a specific template. Users can also submit requests to incorporate custom pre-trained models into the OpenXAI framework by filling a simple form and providing details about model architecture and parameters (See Appendix).
Figure 1: A template of get_explanations function which enables users to incorporate custom explanation methods into OpenXAI.
@@
2) Explainers. OpenXAI provides ready-to-use implementations of six state-of-the-art feature attribution methods, namely, LIME, SHAP, Vanilla Gradients, Gradient x Input, SmoothGrad, and Integrated Gradients. An implementation of a random baseline which randomly assigns importance values to each of the features, and returns these random assignments as explanations is also included. Our implementations of these methods build on other open-source libraries (e.g., Captum [33]) as well as their original implementations. While methods such as LIME and SHAP leverage perturbations of data instances and their corresponding model predictions to learn a local explanation model, they do not require access to the internals of the models or their gradients. On the other hand, Vanilla Gradients, Gradient x Input, SmoothGrad, and Integrated Gradients require access to the gradients of the underlying models but do not need to repeatedly query the models for their predictions (see Table E.2 in Appendix for a brief summary of these methods). These differences influence the efficiency with which explanations can be generated by these methods. OpenXAI provides an abstract Explainer class which enables us to load existing explanation methods as well as integrate new explanation methods.
from openxai import Explainer
exp_method = Explainer(method=‘LIME’)
explanations = exp_method.get_explanations(model, X=inputs, y=labels)
@@
All the explanation methods included in OpenXAI are readily accessible through the Explainer class, and users just have to specify the method name in order to invoke the appropriate method and generate explanations as shown in the above code snippet. Users can easily incorporate their own custom explanation methods into the OpenXAI framework by extending the Explainer class and including the code for their methods in the get_explanations function (see template below) of this class. They can then submit a request to incorporate their custom methods into OpenXAI library by filling a form and providing the GitHub link to their code as well as a summary of their explanation method (See Appendix).
@@
3) Evaluation Metrics. OpenXAI provides implementations and ready-to-use APIs for a set of twenty-two quantitative metrics proposed by prior research to evaluate the faithfulness, stability, and fairness of explanation methods. OpenXAI is the first XAI benchmark to consider all the three aforementioned aspects of explanation reliability. More specifically, we include eight different metrics to measure explanation faithfulness (both with and without ground truth explanations) [34, 48], three different metrics to measure stability [1], and eleven different metrics to measure group-based disparities (unfairness) [12] in the values of the aforementioned faithfulness and stability metrics. The metrics that we choose are drawn from the latest works in explainable AI literature:
@@
a) Ground-truth Faithfulness: Krishna et al. [34] recently proposed a set of evaluation metrics [Feature Agreement (FA), Rank Agreement (RA), Sign Agreement (SA), Signed Rank Agreement (SRA), Rank Correlation (RC), and Pairwise Rank Agreement (PRA)] to capture the similarity between the top-K or a select set of features of any two feature attribution-based explanations. We leverage these metrics to capture the similarity between the explanations output by state-of-the-art methods and the ground-truth explanations constructed using our synthetic data generation process.
@@
b) Predictive Faithfulness: We leverage the metrics outlined by [48, 12] to measure the faithfulness of an explanation when no ground truth is available. This metric, referred to as Prediction Gap on Important feature perturbation (PGI), computes the difference in prediction probability that results from perturbing the features deemed as influential by a given post hoc explanation. Higher values on this metric imply greater explanation faithfulness. We also consider the converse of this metric, Prediction Gap on Unimportant feature perturbation (PGU), which perturbs the unimportant features and measures the change in prediction probability.
@@
c) Stability: We consider the metrics introduced by Alvarez-Melis and Jaakkola [3], Agarwal et al. [1] to measure how robust a given explanation is to small input perturbations. More specifically, we leverage the metrics Relative Input Stability (RIS), Relative Representation Stability (RRS), and Relative Output Stability (ROS) which measure the maximum change in explanation relative to changes in the inputs, model parameters, and output prediction probabilities respectively.

### v5

# OpenXAI: Towards a Transparent Evaluation of Post hoc Model Explanations

@@ Abstract
While several types of post hoc explanation methods have been proposed in recent literature, there is very little work on systematically benchmarking these methods. Here, we introduce OpenXAI, a comprehensive and extensible open-source framework for evaluating and benchmarking post hoc explanation methods. OpenXAI comprises of the following key components: (i) a flexible synthetic data generator and a collection of diverse real-world datasets, pre-trained models, and state-of-the-art feature attribution methods, and (ii) open-source implementations of eleven quantitative metrics for evaluating faithfulness, stability (robustness), and fairness of explanation methods, in turn providing comparisons of several explanation methods across a wide variety of metrics, models, and datasets. OpenXAI is easily extensible, as users can readily evaluate custom explanation methods and incorporate them into our leaderboards. Overall, OpenXAI provides an automated end-to-end pipeline that not only simplifies and standardizes the evaluation of post hoc explanation methods, but also promotes transparency and reproducibility in benchmarking these methods. While the first release of OpenXAI supports only tabular datasets, the explanation methods and metrics that we consider are general enough to be applicable to other data modalities. OpenXAI datasets and models, implementations of state-of-the-art explanation methods and evaluation metrics, are publicly available at this GitHub link.
**footnotetext: Work done by authors when they were at Harvard University.
@@ 1 Introduction
As predictive models are increasingly deployed in critical domains (e.g., healthcare, law, and finance), there has been a growing emphasis on explaining the predictions of these models to decision-makers (e.g., doctors, and judges) so that they can understand the rationale behind model predictions, and determine if and when to rely on these predictions. To this end, various techniques have been proposed in recent literature to generate post hoc explanations of individual predictions made by complex ML models. Several of such local explanation methods output the influence of each of the features on the model’s prediction, and are therefore referred to as local feature attribution methods. Due to their generality, feature attribution methods are increasingly being utilized to explain complex models in medicine, finance, law, and science Elshawi et al. (2019); Ibrahim et al. (2019); Whitmore et al. (2016). Thus, it is critical to ensure that the explanations generated by these methods are reliable so that relevant stakeholders and decision makers are provided with credible information about the underlying models Aivodji et al. (2019).
Prior works have studied several notions of explanation reliability such as faithfulness (or fidelity) Zhou et al. (2021); Liu et al. (2021); Hooker et al. (2018), stability (or robustness) Alvarez-Melis and Jaakkola (2018); Agarwal et al. (2022), and fairness Dai et al. (2022); Balagopalan et al. (2022), and proposed metrics for quantifying these notions. Many of these works also demonstrated through small-scale experiments or qualitative analysis that certain explanation methods are not effective w.r.t. specific notions of reliability. For instance, Alvarez-Melis and Jaakkola (2018) visualized the explanations generated by some of the popular gradient based explanation methods  Simonyan et al. (2014); Shrikumar et al. (2017); Smilkov et al. (2017); Sundararajan et al. (2017) for MNIST images, and showed that they are not robust to small input perturbations. However, it is unclear if such findings generalize beyond the settings studied. More broadly, one of the biggest open questions which has far-reaching implications for the progress of explainable AI (XAI) research is: which explanation methods are effective w.r.t. which notions of reliability and under what conditions?  Krishna et al. (2022). A first step towards answering this question involves systematically benchmarking explanation methods in a reproducible and transparent manner. However, the increasing diversity of explanation methods, and the plethora of evaluation settings and metrics outlined in existing research without standardized open-source implementations make it rather challenging to carry out such benchmarking efforts.
@@
    Our OpenXAI framework currently provides open-source implementations and ready-to-use API interfaces for six state-of-the-art feature attribution methods (LIME, SHAP, Vanilla Gradients, Gradient x Input, SmoothGrad, and Integrated Gradients), and eleven quantitative metrics to evaluate the faithfulness, stability, and fairness of feature attribution methods. In addition, it includes a comprehensive collection of seven real-world datasets spanning diverse real-world domains, and sixteen different pre-trained models. OpenXAI also introduces a novel and flexible synthetic data generator to synthesize datasets of varying sizes, complexity, and dimensionality which facilitate the construction of reliable ground truth explanations (Section 2).
@@
    The OpenXAI framework is easily extensible i.e., researchers and practitioners can readily incorporate custom explanation methods, datasets, predictive models, and evaluation metrics into our framework and benchmarks (Section 2).
@@
    Lastly, using OpenXAI, we perform rigorous empirical benchmarking of the aforementioned state-of-the-art feature attribution methods to determine which methods are effective w.r.t. what notions of reliability across a wide variety of datasets and predictive models (Section 3).
@@
Related Work. Our work builds on the vast literature in explainable AI. Here, we discuss closely related works and their connections to our benchmark (see the Appendix for further discussion).
@@
Evaluation Metrics for Post hoc Explanations: Prior research has studied several notions of explanation reliability, namely, faithfulness (or fidelity), stability (or robustness), and fairness Liu et al. (2021); Zhou et al. (2021); Alvarez-Melis and Jaakkola (2018); Dai et al. (2022). While the faithfulness notion captures how faithfully a given explanation captures the true behavior of the underlying model Zhou et al. (2021); Liu et al. (2021); Hooker et al. (2018), stability ensures that explanations do not change drastically with small perturbations to the input Ghorbani et al. (2019); Alvarez-Melis and Jaakkola (2018). The fairness notion, on the other hand, ensures that there are no group-based disparities in the faithfulness or stability of explanations Dai et al. (2022). To this end, prior works Liu et al. (2021); Slack et al. (2021); Zhou et al. (2021); Alvarez-Melis and Jaakkola (2018); Dai et al. (2022); Hooker et al. (2018) proposed various evaluation metrics to quantify the aforementioned notions. For instance, Petsiuk et al. (2018) measured the change in the probability of the predicted class when important features (as identified by an explanation) are deleted from or introduced into the data instance. A sharp change in the probability implies a high degree of explanation faithfulness. Alvarez-Melis and Jaakkola (2018) loosely quantified stability as the maximum change in the resulting explanations when small perturbations are made to a given instance. Dai et al. (2022) quantified unfairness of explanations as the difference between the faithfulness (or stability) metric values averaged over instances in the majority and the minority subgroups.
XAI Libraries and Benchmarks: Prior works have introduced XAI libraries and benchmarks, the most popular among them being Captum (Kokhlikyan et al., 2020), Quantus (Hedström et al., 2022), XAI-Bench (Liu et al., 2021), and SHAP Benchmark. Below, we provide a brief description for each, and detail how our work differs from them.
While Captum library (Kokhlikyan et al., 2020) is an open-source library that provides implementations and APIs for various state-of-the-art explanation methods, its focus is not on evaluating and/or benchmarking these methods which is the main goal of our work. Quantus library Hedström et al. (2022), on the other hand, provides implementations of certain evaluation metrics to measure the faithfulness and stability/robustness of explanation methods. However, it does not focus on benchmarking explanation methods or providing public dashboards to compare the performance of these methods. Furthermore, the stability/robustness measures (Alvarez-Melis and Jaakkola, 2018) supported by Quantus are somewhat outdated and have been superseded by recently proposed metrics (Agarwal et al., 2022). In addition, Quantus does not support any fairness metrics to evaluate disparities in the quality of explanations which is very important in real-world settings such as healthcare, criminal justice, and policy. In contrast, OpenXAI not only subsumes popular faithfulness and stability/robustness metrics supported by Quantus but also supports 7 new metrics to measure the faithfulness, stability/robustness, as well as the fairness of explanation methods (Agarwal et al., 2022; Dai et al., 2022; Krishna et al., 2022). In addition, OpenXAI focuses on systematically benchmarking state-of-the-art explanation methods and providing public dashboards to readily compare these methods.
SHAP benchmark sha only focuses on evaluating and comparing different variants of SHAP (Lundberg and Lee, 2017a) via certain faithfulness metrics which are similar to the Prediction Gap on Important (PGI) and Unimportant (PGU) feature perturbation metrics outlined in our work. Note that the SHAP benchmark does not include any stability/robustness or fairness metrics. In contrast, OpenXAI benchmarks these metrics extensively across a variety of explanation methods (e.g., LIME, Gradient-based methods).
XAI-Bench Liu et al. (2021) constructed synthetic datasets with ground truth explanations to evaluate the faithfulness of a few explanation methods (e.g., LIME, SHAP, MAPLE). However, recent research argued that their evaluation is unreliable, and predictive models learned using their synthetic datasets may not adhere to the ground truth explanations (Faber et al., 2021). In addition, the aforementioned evaluation is rather limited in scope as synthetic datasets may not even be representative of real-world data (Faber et al., 2021). In contrast, our work not only proposes a novel synthetic data generator that addresses the shortcomings of the synthetic datasets constructed in XAI-Bench but also facilitates the evaluation and benchmarking of the faithfulness, stability, as well as the fairness of 6 state-of-the-art explanation methods on 7 real-world datasets with no ground truth explanations.
In summary, our work makes the following key contributions:
  • We provide implementations and easy-to-use API interfaces for 11 metrics to evaluate the faithfulness, stability, and fairness of explanation methods, 7 of which have not been implemented in prior libraries or benchmarks – the faithfulness metrics Feature Agreement (FA), Rank Agreement (RA), Sign Agreement (SA), Signed Rank Agreement (SRA), Pairwise Rank Agreement (PRA), and the stability metrics Relative Representation Stability (RRS) and Relative Output Stability (ROS).
  • We also introduce a novel and flexible synthetic data generator to synthesize datasets of varying sizes, complexity, and dimensionality to facilitate the construction of reliable ground truth explanations in order to evaluate state-of-the-art explanation methods. Our synthetic data generator addresses the shortcomings of the prior synthetic benchmark (XAI-Bench) by generating synthetic datasets which encapsulate certain key properties, namely, unambiguously defined local neighborhoods, a clear description of feature importances in each local neighborhood, and feature independence. These properties, in turn, allow us to theoretically guarantee that any accurate model trained on our synthetic datasets will adhere to the ground truth explanations of the underlying data.
  • We perform rigorous empirical benchmarking of 6 state-of-the-art feature attribution methods using our OpenXAI framework to determine which methods are effective w.r.t. each of the 11 evaluation metrics across 8 real-world/synthetic datasets, and 16 different predictive models. Note that none of the previously proposed libraries or benchmarks carry out such exhaustive benchmarking efforts across such a wide variety of metrics, models, and datasets.
@@ 2 Overview of OpenXAI Framework
1) Datasets and Predictive Models. The current release of our OpenXAI framework includes a collection of eight different synthetic and real-world datasets. While synthetic datasets allow us to construct ground truth explanations which can then be used to evaluate explanations output by state-of-the-art methods, real-world datasets (where it is typically hard to construct ground truth explanations) help us benchmark these methods in a more realistic manner suitable for practical applications Liu et al. (2021). We would like to note that OpenXAI includes datasets that are widely employed in XAI research to evaluate the efficacy of newly proposed methods and study the behavior of existing methods (Balagopalan et al., 2022; Dai et al., 2022; Dasgupta et al., 2022; Dominguez-Olmedo et al., 2022; Karimi et al., 2020a; Slack et al., 2021; Upadhyay et al., 2021).
Synthetic Datasets: While prior research Liu et al. (2021); Kim et al. (2021) proposed methods to generate synthetic datasets and corresponding ground truth explanations, they all suffer from a significant drawback as demonstrated by Faber et al. (2021) – there is no guarantee that the models trained on these datasets will adhere to the ground truth explanations of the underlying data. This, in turn, implies that evaluating post hoc explanations using the above ground truth explanations would be incorrect since post hoc explanations are supposed to reliably explain the behavior of the underlying model, and not that of the underlying data. To illustrate, let us consider the case where we use aforementioned methods to construct a synthetic dataset with features $A,B,$ $C$, and $D$ such that the ground truth labels only depend on features $A$ and $B$ i.e., the ground truth explanation of the underlying data indicates that features $A$ and $B$ are most important. If we train a model on this data and if features $A$ and $B$ are correlated with $C$ and $D$ respectively, then the resulting model may base its predictions on $C$ and $D$ (and not $A$ and $B$) and still be very accurate. If a post hoc explanation of this model then (correctly) indicates that the most important features of the model are $C$ and $D$, this explanation may be deemed incorrect if we compare it against the ground truth explanation of the underlying data. This problem further exacerbates as we increase the complexity of the ground truth labeling function Faber et al. (2021).
@@
Real-world Datasets: In the current release of OpenXAI, we include seven real-world datasets that are highly diverse in terms of several key properties. They comprise of data spanning multiple real-world domains (e.g., finance, lending, healthcare, and criminal justice), varying dataset sizes (e.g., small vs. large-scale), dimensionalities (e.g., low vs. high dimensional), class imbalance ratios, and feature types (e.g., continuous vs. discrete). We focus on tabular data in this release as such data is commonly encountered in real-world applications where explainability is critical Verma et al. (2020), and has also been widely studied in XAI literature Liu et al. (2021). Table 1 provides a summary of the real-world datasets currently included in OpenXAI. See Section D.1 in the Appendix for detailed descriptions of individual datasets. While these real-world datasets are primarily drawn from prior research and existing repositories, OpenXAI provides comprehensive data loading and pre-processing capabilities to make these datasets XAI-ready (further details below).
Dataset
Size
# Features
Feature Types
Feature Information
Balanced
Synthetic Data
5,000
20
continuous
synthetic
German Credit Dua and Graff (2017)
1,000
20*
discrete, continuous
demographic, personal, financial
HELOC FICO (2022)
9,871
23
continuous
demographic, financial
Adult Income Yeh and Lien (2009)
45,222
13
discrete, continuous
demographic, personal, education/employment, financial
COMPAS Jordan and Freiburger (2015)
6,172
7
discrete, continuous
demographic, personal, criminal
Give Me Some Credit Freshcorn (2022)
102,209
10
discrete, continuous
demographic, personal, financial
Pima-Indians Diabetes Smith et al. (1988)
768
8
discrete, continuous
demographic, healthcare
Framingham Heart Study Fra
3,658
15
continuous
demographic, healthcare
*expands to 60 features after one hot encoding of discrete features
Dataloaders: OpenXAI provides a ReturnLoaders function that can be used to load the aforementioned collection of synthetic and real-world datasets as well as any other custom datasets, and ensures that they are XAI-ready. More specifically, this class takes as input the name of an existing OpenXAI dataset and outputs a train set which can then be used to train a predictive model, and a test set which can be used to generate local explanations of the trained model. The code snippet below shows how to import the ReturnLoaders function and load an existing OpenXAI dataset. from openxai.dataloader import ReturnLoaders
trainloader, testloader = ReturnLoaders(data_name=‘german’, download=True)
inputs, labels = next(iter(testloader))
Pre-trained models: We train two classes of predictive models (artificial neural networks and logistic regression models) and incorporate them into the OpenXAI framework so that they can be readily used for benchmarking explanation methods. The code snippet below shows how to load OpenXAI’s pre-trained models using our LoadModel function. from openxai.model import LoadModel
@@
2) Explainers. OpenXAI provides ready-to-use implementations of six state-of-the-art feature attribution methods, namely, LIME, SHAP, Vanilla Gradients, Gradient x Input, SmoothGrad, and Integrated Gradients. An implementation of a random baseline which randomly assigns importance values to each of the features, and returns these random assignments as explanations is also included. Our implementations of these methods build on other open-source libraries (e.g., Captum Kokhlikyan et al. (2020)) as well as their original implementations. While methods such as LIME and SHAP leverage perturbations of data instances and their corresponding model predictions to learn a local explanation model, they do not require access to the internals of the models or their gradients. On the other hand, Vanilla Gradients, Gradient x Input, SmoothGrad, and Integrated Gradients require access to the gradients of the underlying models but do not need to repeatedly query the models for their predictions (see Table D.2 in Appendix for a brief summary of these methods). These differences influence the efficiency with which explanations can be generated by these methods. OpenXAI provides an abstract Explainer class which enables us to load existing explanation methods as well as integrate new explanation methods. from openxai.explainer import Explainer
exp_method = Explainer(method=‘shap’, model=model)
explanations = exp_method.get_explanations(inputs)
@@
All the explanation methods included in OpenXAI are readily accessible through the Explainer class, and users just have to specify the method name in order to invoke the appropriate method and generate explanations as shown in the above code snippet. Users can easily incorporate their own custom explanation methods into the OpenXAI framework by extending the Explainer class and including the code for their methods in the get_explanations function (see template above) of this class. They can then submit a request to incorporate their custom methods into OpenXAI.
@@
3) Evaluation Metrics. OpenXAI provides implementations and ready-to-use APIs for a set of twenty-two quantitative metrics proposed by prior research to evaluate the faithfulness, stability, and fairness of explanation methods. OpenXAI is the first XAI benchmark to consider all the three aforementioned aspects of explanation reliability. More specifically, we include eight different metrics to measure explanation faithfulness (both with and without ground truth explanations) Krishna et al. (2022); Petsiuk et al. (2018), and three different metrics to measure stability Agarwal et al. (2022). Below, we briefly describe these metrics. Detailed descriptions of all the metrics along with notation and equations are included in the Appendix.
@@
a) Ground-truth Faithfulness:  Krishna et al. (2022) recently proposed six evaluation metrics to capture the similarity between the top-K or a select set of features of any two feature attribution-based explanations. We leverage these metrics to capture the similarity between the explanations output by state-of-the-art methods and the ground-truth explanations constructed using our synthetic data generation process. These metrics and their definitions are given as follows: i) Feature Agreement (FA) which computes the fraction of top-K features that are common between a given post hoc explanation and the corresponding ground truth explanation, ii) Rank Agreement (RA) metric which measures the fraction of top-K features that are not only common between a given post hoc explanation and the corresponding ground truth explanation, but also have the same position in the respective rank orders, iii) Sign Agreement (SA) metric which computes the fraction of top-K features that are not only common between a given post hoc explanation and the corresponding ground truth explanation, but also share the same sign (direction of contribution) in both the explanations, iv) Signed Rank Agreement (SRA) metric which computes the fraction of top-K features that are not only common between a given post hoc explanation and the corresponding ground truth explanation, but also share the same feature attribution sign (direction of contribution) and position (rank) in both the explanations, v) Rank Correlation (RC) metric which computes Spearman’s rank correlation coefficient to measure the agreement between feature rankings provided by a given post hoc explanation and the corresponding ground truth explanation, and vi) Pairwise Rank Agreement (PRA) metric which captures if the relative ordering of every pair of features is the same for a given post hoc explanation as well as the corresponding ground-truth explanation.
@@
b) Predictive Faithfulness: We leverage the metrics outlined by Petsiuk et al. (2018); Dai et al. (2022) to measure the faithfulness of an explanation when no ground truth is available. This metric, referred to as Prediction Gap on Important feature perturbation (PGI), computes the difference in prediction probability that results from perturbing the features deemed as influential by a given post hoc explanation. Higher values on this metric imply greater explanation faithfulness. We also consider the converse of this metric, Prediction Gap on Unimportant feature perturbation (PGU), which perturbs the unimportant features and measures the change in prediction probability.
@@
c) Stability: We consider the metrics introduced by Alvarez-Melis and Jaakkola (2018); Agarwal et al. (2022) to measure how robust a given explanation is to small input perturbations. More specifically, we leverage the metrics Relative Input Stability (RIS), Relative Representation Stability (RRS), and Relative Output Stability (ROS) which measure the maximum change in explanation relative to changes in the inputs, model parameters, and output prediction probabilities respectively.